# The Multiband Imaging Survey for High-Alpha PlanetS (MISHAPS) I: Preliminary Constraints on the Occurrence Rate of Hot Jupiters in 47 Tucanae

### v1

@@ 8 Discussion
We have not reproduced Masuda & Winn (2017)’s methodology, but if we assume a similar two thirds detection efficiency for an appropriately weighted distribution, our limit combined with Masuda & Winn (2017) would yield a combined $N_{1}=1776$ and an occurrence rate limit of $f_{\rm HJ}<0.17$%, which would just about rule out MW17’s prediction of $f_{\rm HJ}=0.18$% that assumes no metallicity dependence.
If we assume that the strong dependence of giant planet occurrence rate on metallicity holds for globular cluster planet formation, then our new limits do not place particularly interesting constraints on 47 Tuc’s hot Jupiter occurrence rate. Johnson et al. (2010) fit for a giant planet occurrence rate that depends on host mass and metallicity of the form
@@
finding $\alpha=1.0\pm 0.3$ and $\beta=1.2\pm 0.2$ for giant planets with semimajor axis $<2.5$ AU. Similar estimates for hot Jupiters yield an even stronger metallicity dependence, with Guo et al. (2017) finding a metallicity dependence of $f_{\rm HJ}\propto 10^{2.1\pm 0.7[{\rm Fe}/{\rm H}]}$, and Petigura et al. (2018) found $f_{\rm HJ}\propto 10^{3.4_{-0.8}^{+0.9}[{\rm Fe}/{\rm H}]}$, both studies having relatively small numbers of hot Jupiters and do not account for a possible host mass dependence. If we are optimistic and adopt the Johnson et al. (2010) metallicity relation and apply it to 47 Tuc’s metallicity $[{\rm Fe}/{\rm H}]=-0.78$ dex (Cordero et al., 2014) and use MW17’s $f_{\rm HJ}=0.24$% occurrence rate for hot Jupiters with $P<10$ days, we can predict the hot Jupiter occurrence rate in 47 Tuc to be $0.028$%. However, if we can assume that $\alpha$-elements actually set the planet occurrence rate and not iron, then by replacing [Fe/H] with $[\alpha/{\rm H}]=[{\rm Fe}/{\rm H}]+[\alpha/{\rm Fe}]$ in equation 25, we can predict $f_{\rm HJ}=0.055$% by adopting $[\alpha/{\rm Fe}]=0.3$ dex  (Cordero et al., 2014). While our occurrence limits apply to slightly different planet radius and period ranges, the ones that combine our search with G00 are within a factor of 22223333 of the $\alpha$ prediction. A 95% upper limit that rules out the $\alpha$ prediction will require a survey sensitivity of $N_{1}=5450$.
[…]
@@ 9 Conclusions
Though we find no planets in 47 Tuc, we are able to constrain the giant planet occurrence rate to an upper limit of $<$0.43% at 95% confidence. When combined with the results of Gilliland et al. (2000), we can place an even tighter limit of $<$0.11%. We present detailed analysis of 39 interesting targets, consisting of 4 eclipsing binaries and 35 transit candidates. Using our two-step vetting process, we are able to reject all 35 of these transit candidates as false positives without follow-up observations.
@@ Acknowledgments
Based on observations at Cerro Tololo Inter-American Observatory, NSF’s NOIRLab (NOIRLab Prop. IDs 2019A-0315, PI: M. Penny; 2022B-511337, PI: A. Crisp), which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation.
Raw imaging data were processed with the DECam Community Pipeline (Valdes et al., 2014).
@@ Appendix B Additional Candidate Information
The following plots show the lightcurves and images for each of our 35 non-EB candidates. For each candidate, the left panel is the lightcurve of a night with a detection, the middle panel is the in-transit difference image stack, and the right panel is the reference image. For candidates with multiple detections, there is one row of plots per detection. In the lightcurve panels, the $r-$band data are plotted in blue, the $z-$band data are plotted in red, and the boxcar models from the search are plotted in green.

### v2

@@ 8 Discussion
We have not reproduced K. Masuda & J. N. Winn (2017)’s methodology, but if we assume a similar two thirds detection efficiency for an appropriately weighted distribution, our limit combined with K. Masuda & J. N. Winn (2017) would yield a combined $N_{1}=1776$ and an occurrence rate limit of $f_{\rm HJ}<0.17$%, which would just about rule out MW17’s prediction of $f_{\rm HJ}=0.18$% that assumes no metallicity dependence.
If we assume that the strong dependence of giant planet occurrence rate on metallicity holds for globular cluster planet formation, then our new limits do not place particularly interesting constraints on 47 Tuc’s hot Jupiter occurrence rate. J. A. Johnson et al. (2010) fit for a giant planet occurrence rate that depends on host mass and metallicity of the form
@@
finding $\alpha=1.0\pm 0.3$ and $\beta=1.2\pm 0.2$ for giant planets with semimajor axis <2.5<2.5 AU. Similar estimates for hot Jupiters yield an even stronger metallicity dependence, with X. Guo et al. (2017) finding a metallicity dependence of $f_{\rm HJ}\propto 10^{2.1\pm 0.7[{\rm Fe}/{\rm H}]}$, and E. A. Petigura et al. (2018) found $f_{\rm HJ}\propto 10^{3.4_{-0.8}^{+0.9}[{\rm Fe}/{\rm H}]}$, both studies having relatively small numbers of hot Jupiters and do not account for a possible host mass dependence. If we are optimistic and adopt the J. A. Johnson et al. (2010) metallicity relation and apply it to 47 Tuc’s metallicity ${\rm Fe}/{\rm H}]=-0.78$ dex (M. J. Cordero et al., 2014) and use MW17’s $f_{\rm HJ}=0.24$% occurrence rate for hot Jupiters with $P<10$ days, we can predict the hot Jupiter occurrence rate in 47 Tuc to be 0.0280.028%. However, if we can assume that $\alpha$-elements actually set the planet occurrence rate and not iron, then by replacing [Fe/H] with $\alpha/{\rm H}]=[{\rm Fe}/{\rm H}]+[\alpha/{\rm Fe}]$ in equation 26, we can predict $f_{\rm HJ}=0.055$% by adopting $\alpha/{\rm Fe}]=0.3$ dex  (M. J. Cordero et al., 2014). While our occurrence limits apply to slightly different planet radius and period ranges, the ones that combine our search with G00 are within a factor of 2233 of the $\alpha$ prediction. A 95% upper limit that rules out the $\alpha$ prediction will require a survey sensitivity of $N_{1}=5450$.
This work is only the first step toward our goal of providing improved occurrence rate limits in 47 Tuc. It is already 50% more sensitive than the previous ground-based search of 47 Tuc (W05), and per star our sensitivity is within 10% of G00’s. We note that our analysis is the first globular cluster planet search to quantify the human vetting detection efficiency, which we find is high (${\sim}90$%) but significantly lower than 1, especially at longer periods. However, we have only determined this efficiency for the first vetting step; it is likely that the remaining vetting steps we take also are not 100% efficient. Additionally, our analysis is the first 47 Tuc transit search to exclude Milky Way field stars and stars in the SMC via proper motion cuts, which provide a small but significant contribution to the stellar sample in the magnitude range $r\approx 18$2020 due to the 47 Tuc main sequence crossing the SMC giant branch in the regions of the red clump and red giant branch bump. In our next paper, we will improve our occurrence rate limits by incorporating the central cluster chips into our analysis, including faint stars missed by Gaia but on which we can effectively detect transits, by upgrading our search to incorporate more realistic transit shapes, and to include data from nights that were not as bad as originally assumed, or that were gathered after the inclusion cutoff for this paper. Combined, we anticipate these additions to increase $N_{1}$ to the point where we can challenge the most optimistic predictions of planet occurrence rate if we again include G00.
@@ 9 Conclusions
Though we find no planets in 47 Tuc, we are able to constrain the giant planet occurrence rate to an upper limit of <<0.43% at 95% confidence. When combined with the results of R. L. Gilliland et al. (2000), we can place an even tighter limit of <<0.11%. We present detailed analysis of 39 interesting targets, consisting of 4 eclipsing binaries and 35 transit candidates. Using our two-step vetting process, we are able to reject all 35 of these transit candidates as false positives without follow-up observations.
@@ Acknowledgments
We would also like to thank the referee for their helpful comments and suggestions.
@@
Based on observations at Cerro Tololo Inter-American Observatory, NSF’s NOIRLab (NOIRLab Prop. IDs 2019A-0315, PI: M. Penny; 2022B-511337, PI: A. Crisp). This research also uses services or data provided by the Astro Data Lab, which is part of the Community Science and Data Center (CSDC) Program of NSF NOIRLab. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA), Inc. under a cooperative agreement with the U.S. National Science Foundation.
Raw imaging data were processed with the DECam Community Pipeline (F. Valdes et al., 2014).
@@ Appendix B Additional Candidate Information
The following plots show the lightcurves and images for each of our 35 non-EB candidates. In each plot, the left panel is the lightcurve of a night with a detection, the middle panel is the in-transit difference image stack, and the right panel is the reference image. In the lightcurve panels, the $r$-band data are plotted in blue, the $z$-band data are plotted in red, the boxcar model from the search is plotted in solid green, and the median magnitude is plotted in dashed green. The values displayed on the plot are — clockwise from top left — the estimated stellar radius, estimated planet radius, transit duration, transit depth, and S/N of the detection. For candidates with multiple detections, there is one row of plots per detection.